# Activation Function Design Sustains Plasticity in Continual Learning

### v1

@@ 7 Continual Reinforcement Learning
Therefore, to investigate plasticity loss in a continual RL setting, we train a single PPO agent Schulman et al. (2017) on a fixed and repeating sequence of four MuJoCo locomotion tasks using Gymnasium Towers et al. (2024): HalfCheetah-v4 $\rightarrow$ Hopper-v4 $\rightarrow$ Walker2d-v4 $\rightarrow$ Ant-v4 $\rightarrow$ HalfCheetah-v4 $\rightarrow$ Hopper-v4 $\rightarrow$ $\cdots$. The agent cycles through this sequence three times, training for 1M timesteps per environment per cycle (total 12M). Episodes terminate early on invalid configurations following Dohare et al. (2024) (e.g., falls for Hopper/Walker2d, unstable heights for Ant); HalfCheetah runs full-length. The policy and value networks share a two-layer MLP backbone (256 units each) that is updated across all tasks. For each environment we attach a lightweight input adapter (to map its observation space) and a task-specific output head (for its action space); both persist across cycles and continue learning when the environment reappears. Hyperparameters are in App. LABEL:tab:hyper_sweep.
Metric Swish PReLU Sigmoid Rand. Smooth-Leaky
Mean $\pm$ 95% CI 1120.28  $\pm$ 229.22 684.88  $\pm$ 115.90 993.61  $\pm$ 190.38 1310.94  $\pm$ 98.37
Table 3: Mean Plasticity Score across 5 seeds (higher is better). We only report the top-performing activations. See Table F1 for a full comparison of all activations. Values are reported as mean $\pm$ 95% CI half-width. For bootstrap CIs, which can be asymmetric, we report the larger one-sided margin so the printed $\pm$ interval conservatively covers the equal-tailed 95% CI [L,U].
Figure 4: Plasticity Score across 5 seeds (95% bootstrap CIs) showing a complete sequence of 3 cycles across all 4 environments. The table 3 reports the mean of this score across seeds, only showing the top-performing activations to avoid clutter.
@@
For each run $r$: (i) in the last cycle of training, compute the mean episodic return for each environment over the final $p{=}15\%$ of that environment’s cycle (steady-state); (ii) take the median across environments to avoid any single task dominating the summary. This per-run statistic is the Plasticity Score reported in Table 3. This aggregates performance across multiple environments while emphasizing late-cycle adaptability.
Across seeds, Rand. Smooth-Leaky achieves the highest plasticity score, outperforming Swish, Sigmoid, and PReLU. These results highlight that smooth and randomized variants provide more robust adaptation and retention across sequential environments compared to classical activations. Figure 4 complements this analysis by showing reward trajectories across cycles for the top-performing activations.
@@
Anchoring plasticity loss to a single mechanism risks conflating cause and effect. Prior work (Berariu et al., 2021) distinguishes trainability—reduced ability to lower loss on new data (Dohare et al., 2021; Elsayed & Mahmood, 2024; Lyle et al., 2022)—from generalizability—reduced performance on unseen data (Ash & Adams, 2020; Zilly et al., 2021). Mitigating plasticity loss should improve both, not merely memorize the latest data Lee et al. (2024); nevertheless, their relationship remains unsettled, and terminology is often mixed Klein et al. (2024). In brief: loss of trainability (no effective parameter updates) can cause loss of functional plasticity, but the reverse need not hold—models may still update yet fail to transfer gains.
@@
Thus, we can argue that loss of plasticity loss has two faces in RL: (i) train-side adaptation—can the agent still improve on the data it now collects?—and (ii) transfer to perturbed test conditions. We report one metric for each. Plasticity Score (Sec. 7, Tab. 3) summarizes late-cycle train-side performance. Generalization Gap compares the same policy on matched test variants (perturbed dynamics/seeds; App. F): for each cycle $c$ and environment $e$, $\mathrm{GAP}_{c,e}=R^{\text{train}}_{c,e}-R^{\text{test}}_{c,e}$, where $R$ is the expected return (measured at the end of cycle $c$). Thus $\Delta(\mathrm{GAP}_{e})=\mathrm{GAP}_{3,e}-\mathrm{GAP}_{1,e}$; larger $\Delta$ means the train–test gap widened over time (worse transfer). We aggregate $\Delta$ per activation by the median across environments.
Among the high-plasticity activations, Rand. Smooth-Leaky stands out, while not uniformly best per environment: it attains the lowest median $\Delta(\mathrm{GAP}_{\text{e}})$ across MuJoCo (Tab. 4), indicating the least widening of the train–test gap on average—and on Humanoid-v5 it even narrows the gap ($\Delta<0$). In contrast, Swish and Sigmoid score highly on late-cycle Plasticity Score (Fig. 4, Tab. 3) yet exhibit positive $\Delta(\mathrm{GAP}_{\text{e}})$, revealing a trainability–transfer tension. PReLU is competitive on some environments (best $\Delta$ on Ant-v5 and Hopper-v5) but has a higher median overall.
@@
These metrics are complementary, not redundant. Plasticity Score asks “did the agent remain adaptable on the data it collected?”, while $\Delta(\mathrm{GAP}_{\text{e}})$ asks “did that adaptation carry to perturbed tests?”. Our activation designs primarily target sustained train-side plasticity (strict floor, moderate leak, $C^{1}$ transition)—a prerequisite for learning under shift. Improving transfer likely requires coupling activations with policy/regularization choices (e.g., band-regularizing the learned leak toward target bounds, entropy/KL tuning, dynamics randomization). We report both metrics to surface, not mask, this open question.
Activation HalfCheetah-v5 Humanoid-v5 Ant-v5 Hopper-v5 median $\Delta$
Sigmoid 1521.45 18.92 276.48 152.14 214.31
PReLU 839.60 128132.00 94.17 -22.35 466.88
Swish (SiLU) 627.24 -127118.70 1035.22 20.88 324.06
Rand. Smooth-Leaky 103.80 -1236645.00 563.09 19.22 61.51
Table 4: Change in Generalization Gap from cycle 1 to cycle 3, per environment and activation (rounded to 2 decimals). $\Delta(\mathrm{GAP}_{\text{e}})=\mathrm{GAP}_{\text{3,e}}-\mathrm{GAP}_{\text{1,e}}$; $\Delta(\mathrm{GAP}_{\text{e}})>0$ means the train–test gap widened (worse transfer). The rightmost column is the median across environments to reduce sensitivity to scale differences. Full results in Table F2. We interpret large negative generalization values in App.F.1.
@@ Appendix A Characterization of Activation Function Properties
ReLU Nair & Hinton (2010)
@@
GeLU Hendrycks & Gimpel (2016)
@@
*Gradients are small but non‑zero except at extreme inputs (risk of effective inactivity via saturation). ELU is $C^{1}$ only for $\alpha=1$. SELU has a small derivative jump at $x=0$ because $\alpha\neq 1$. SELU’s self-normalizing behavior holds under the prescribed $(\lambda,\alpha)$. Unconstrained activations are smooth, non-monotonic, and non-saturating; outputs can grow large under training. Rational do not have “slopes” like Leaky-ReLU/PReLU; they learn polynomial coefficients. Each branch retains the ReLU dead zone, but concatenation ensures that for any input at least one branch is active, avoiding global inactivity. Proposed in this work. Potentially non-monotonic depending on the choice of slope parameter (and on the randomization bounds for R-Smooth-Leaky).
@@ Appendix B Experimental and Architectural Design
B.1 Limitations and implications of multi-parameter activation design.
Smooth-Leaky and Randomized Smooth-Leaky expose several shape-controlling hyperparameters. This enlarges the tuning space—and thus the cost—but also grants useful control, as reflected in our benchmarks. A common response is to replace sweeps with adaptive or learnable parameters. However, as shown in Section 3, adaptive, granular learnable slopes (e.g., PReLU) underperformed in all our settings, while bounded, stochastic constraints (e.g., RReLU) worked better. In continual learning, where short-term objectives dominate, unconstrained learned parameters can drift to task-local optima that hurt performance across tasks. Attempts to automate shape learning (Bingham & Miikkulainen, 2023; Bingham et al., 2020; Manessi & Rozza, 2018) with highly expressive forms (e.g., Rational activations Delfosse et al. (2020)) also lagged behind our first-principles designs here. This suggests that we have not yet reconciled expressivity with robust automation. A promising direction is to rethink where and on what timescale hyperparameters are adapted—potentially decoupling their updates from the main training loop to better handle non-stationarity.
B.2 CReLU integration.
@@
B.3 Hyper-parameter Sweeps for Activation Functions
@@
Leaky ReLU $\alpha$ 0.01, 0.05, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0
ELU / CELU $\alpha$ 0.1, 0.5, 1.0, 1.5, 2.0, 2.2, 2.4, 2.6, 2.8, 3.0, 3.3, 3.5, 3.6, 3.9
SELU $\alpha$ 1.0, 1.3, 1.673, 2.0, 2.3, 2.4, 2.6, 2.8, 3.0, 3.1, 3.3, 3.5, 3.7
Swish / GeLU $\beta$ 0.01, 0.05, 0.1, 0.5, 0.8, 1.0
@@
B.4 i.i.d. vs Class-Incremental Continual Learning comparison Hyperparameters
@@
B.5 Understanding Class-Incremental Continual Learning Metrics
To properly understand the choice of metrics throughout this paper, we first need to define and explain some of the classical metrics used in continual learning—Final Average Accuracy (ACC), Backward Transfer (BWT), and Forward Transfer (FWT)—and relate them to loss of plasticity. Consider a learning process unfolding over N+1N{+}1 sequential phases/tasks $t\!=\!0,\dots,N$. Between consecutive phases t1t{-}1 and $t$ ($t\!\geq\!1$), the data distribution, environment dynamics, reward signal, or task objective may change, inducing non-stationarity; we denote the distribution at phase $t$ by $E_{t}$. The model enters phase $t$ with parameters $\theta_{t-1}$ and is updated for $T_{\text{phase}}$ adaptation steps333“Steps” may be epochs, mini-batches, or timesteps, depending on the domain., producing intermediate iterates $\theta_{t,k}$ ($k\!=\!1,\dots,T_{\text{phase}}$) and terminating at $\theta_{t}\!=\!\theta_{t,T_{\text{phase}}}$.
@@ B.5 Understanding Class-Incremental Continual Learning Metrics
B.5.1 Different Accuracy metrics across Continual Learning
@@ Appendix C Expanding on Case Study 1: Negative‑Slope ‘Goldilocks Zone’
C.1 ‘Goldilocks zone’ for negative slopes (with shape-matched comparison)
Matching each activation’s effective negative slope to that of Leaky-ReLU shows that—across all tested shape-controlling parameters—no alternative attains equal or higher final average accuracy, reinforcing the presence of a ‘Goldilocks zone’. Consistently, Fig. C1b shows the fraction of dead units declining as the effective slope enters the green band, aligning with the view that dead units contribute to plasticity loss; accuracy peaks in the same region.
Figure C1: Left: Final accuracy vs. effective negative slope $\bar{s}$. Right: Dead-unit fraction vs. $\bar{s}$. Linear-leak families (blue) peak for $\bar{s}\!\in\![0.6,0.9]$. Smooth-tailed activations (ELU/CELU/SELU, Swish, GELU) are plotted on the same $\bar{s}$ axis; they underperform within the ‘Goldilocks zone’ and only approach the linear-leak peak when $\bar{s}\!>\!1$, reflecting concentrated near-zero responsiveness and vanishing tails.
[…]
On the other hand, we also check their $\lambda_{\max}$ curves (Fig. C2, which also sit noticeably above Leaky-ReLU’s optimal low-curvature trough. This suggests that the non-constant, often exponential or logistic, shape of their negative tails fails to flatten landscape curvature as effectively as Leaky-ReLU’s constant slope, even when average gradient throughput ($\bar{s}$) is matched. However, at extreme effective slopes ($\bar{s}>1.0$) where Leaky-ReLU’s performance collapses due to curvature explosion, functions like ELU and SELU maintain more moderate curvature and salvage higher ACC (around 22%), indicating better robustness at this specific extreme.
C.2 Failure Modes for Slope Magnitude.
Continuing with the explanation in Sec. 3; the so-called ‘Goldilocks phenomenon’ highlights two distinct failure modes at opposite ends of the negative slope range. As shown in Fig. C2 (left) the largest Hessian eigenvalue, $\lambda_{\max}$, follows a U-shaped curve. Curvature is high when the negative slope $\bar{s}\approx 0$ (due to many "dead" units) and rises again when $\bar{s}>0.9$ (as the function behaves like an identity map). The ‘Goldilocks zone’ lies between $\bar{s}\approx 0.6$ and $0.9$, where $\lambda_{\max}$ is at its lowest. The effective rank, illustrated in Fig. C2 (right) offers more insight. The rank is lowest at both extremes: near $\bar{s}\approx 0$ because of widespread unit death, and at very high $\bar{s}$ where extreme curvature creates a stiff, low-dimensional learning subspace.
Optimal plasticity is achieved in the ‘Goldilocks zone’ where the network balances having active learning directions with a compact representation and flat curvature. Therefore, loss of plasticity is caused by unit death at very small negative slopes and by high curvature at very large negative slopes. A ‘Goldilocks zone’, independently of the range that it reports, is currently found empirically. However, its mere existence suggests that there are optimal configurations.
Figure C2: Left: effective rank of the gradient Gram matrix. Right: dominant Hessian eigenvalue $\lambda_{\max}$. Smooth-tailed activations (swish, celu) show spikes at large $\bar{s}$, while constant-slope leaks remain comparatively stable.
We also investigated the primary drivers of functional plasticity loss, we compare across a set of representative activation configurations and all negative slope variations. As provided in the main section results, the correlation between Final Average Accuracy (ACC) and dead unit fraction is strong and significant (Pearson (r)=0.59(r)=-0.59 ($p{=}5.7{\times}10^{-8}$), as seen in Fig. C3, right. In contrast, the correlation between ACC and final scaled gradient norms is weak and not significant (Pearson (r)=0.28(r){=}\,-0.28) (p=0.019p{=}0.019) (Fig. C3, left).
Settings that maintain dead unit fraction below approximately 8% consistently achieve high ACC (27–33%), whereas dead unit rates above 25% lead to ACC collapsing to 20% or lower. This indicates that while gradient magnitude is a factor, the rise in dead units (gradient starvation) is a more direct and reliable indicator of plasticity loss than the raw magnitude of gradients, which can be small in well-adapted, flat minima or large in failing networks with high curvature. The results underscore that plasticity is not about maximizing any single statistic but about achieving a delicate balance, primarily tuned by the negative slope characteristics, to ensure unit survival, maintain diverse learning directions, and promote a benign loss landscape.
@@
C.3 Adaptive, granular slopes are useful—but need guidance to stay in-band

### v2

@@ 7 Continual Reinforcement Learning
Therefore, to investigate plasticity loss in a continual RL setting, we train a single PPO agent Schulman et al. (2017) on a fixed and repeating sequence of four MuJoCo locomotion tasks using Gymnasium Towers et al. (2024): HalfCheetah-v4 $\rightarrow$ Hopper-v4 $\rightarrow$ Walker2d-v4 $\rightarrow$ Ant-v4 $\rightarrow$ HalfCheetah-v4 $\rightarrow$ Hopper-v4 $\rightarrow$ $\cdots$. The agent cycles through this sequence three times, training for 1M timesteps per environment per cycle (total 12M). Episodes terminate early on invalid configurations following Dohare et al. (2024) (e.g., falls for Hopper/Walker2d, unstable heights for Ant); HalfCheetah runs full-length. The policy and value networks share a two-layer MLP backbone (256 units each) that is updated across all tasks. For each environment we attach a lightweight input adapter (to map its observation space) and a task-specific output head (for its action space); both persist across cycles and continue learning when the environment reappears. We trained using the Adam optimizer. Hyperparameters are in App. LABEL:tab:hyper_sweep.
Metric Swish PReLU Sigmoid Rand. Smooth-Leaky Smooth-Leaky
IQM $\pm$ 95% CI 0.3149  $\pm$ 0.071 0.2716  $\pm$ 0.038 0.3329  $\pm$ 0.059 0.3875  $\pm$ 0.038 0.3305  $\pm$ 0.037
Table 3: Average Plasticity Score across 5 seeds (higher is better). We only report the top-performing activations. See Table F1 for a full comparison of all activations. Values are reported as Min-Max Normalized IQM mean $\pm$ 95% CI half-width.
Figure 6: Plasticity Score across 5 seeds (95% bootstrap CIs) showing a complete sequence of 3 cycles across all 4 environments. The table 3 reports the Min-Max Normalized IQM of this score across seeds, only showing the top-performing activations to avoid clutter.
@@
To rigorously aggregate performance across environments with distinct reward scales, we employ a normalized scoring protocol following Agarwal et al. (2021). For each run $r$ we compute the mean episodic return of the last cycle’s steady-state (final $15\%$ of steps). This value $R_{r,e}$ for environment $e$ is Min-Max normalized: $S_{r,e}=\frac{R_{r,e}-R_{e}^{\min}}{R_{e}^{\max}-R_{e}^{\min}}$. Global bounds $R_{e}^{\min},R_{e}^{\max}]$ are determined via a robust percentile sweep across all activations. To prevent physics simulation instabilities in failure cases (e.g., Humanoid-v5) from skewing the mean with unbounded negative values, we apply a stability floor: rewards below a functional failure threshold are clipped to the lower bound ($S_{r,e}=0$). Finally, we report the Interquartile Mean (IQM) to provide a robust measure of expected plasticity. By filtering outliers while accounting for performance magnitude, the resulting Plasticity Score $\in[0,1]$ reflects the agent’s efficiency relative to the environment’s theoretical maximum.
@@
Anchoring plasticity loss to a single mechanism risks conflating cause and effect. We distinguish trainability (learning capacity on current data) from generalizability (transfer to unseen variations). Prior work (Berariu et al., 2021) distinguishes trainability—reduced ability to lower loss on new data (Dohare et al., 2021; Elsayed and Mahmood, 2024; Lyle et al., 2022)—from generalizability—reduced performance on unseen data (Ash and Adams, 2020; Zilly et al., 2021). Mitigating plasticity loss should improve both, not merely memorize the latest data Lee et al. (2024); nevertheless, their relationship remains unsettled, and terminology is often mixed Klein et al. (2024). In brief: loss of trainability (no effective parameter updates) can cause loss of functional plasticity, but the reverse need not hold—models may still update yet fail to transfer gains.
@@
This failure of plasticity has two faces in RL: (i) train-side adaptation—can the agent still improve on the data it now collects?—and (ii) transfer to perturbed test conditions (e.g., randomized MuJoCo friction Dohare et al. (2024)). We report two complementary metrics. The Plasticity Score (IQM) summarizes late-cycle trainability. The Generalization Gap ($\Delta\mathrm{GAP}$) measures how much the train–test gap widens over time (App. F). For each cycle $c$ and environment $e$, $\mathrm{GAP}_{c,e}=R^{\text{train}}_{c,e}-R^{\text{test}}_{c,e}$, where $R$ is the expected return (measured at the end of cycle $c$). Thus $\Delta(\mathrm{GAP}_{e})=\mathrm{GAP}_{3,e}-\mathrm{GAP}_{1,e}$; larger $\Delta$ means the train–test gap widened over time (worse transfer). We aggregate $\Delta$ per activation using the Interquartile Mean (IQM) across environments. Crucially, these metrics must be analyzed together.
A high Plasticity Score is not preferable if it comes with a large, widening gap, just as a small gap is meaningless if absolute rewards are near zero (as seen in physics failures). Our two-metric system captures this tension. While Rand. Smooth-Leaky suffers from stability issues in Humanoid (zero rewards and thus a zero gap), in stable environments like Ant and Cheetah where it maximizes the Plasticity Score, it also attains a lower IQM $\Delta(\mathrm{GAP}_{\text{e}})$ compared to baselines (Tab. 4). This indicates that its trainability does not come at the cost of overfitting; rather, it encourages solutions that transfer well, provided the dynamics remain stable. In contrast, Sigmoid achieves a moderate Plasticity Score via safety—its bounded nature prevents divergence in Humanoid—yet it exhibits a larger widening of the gap in other tasks.
@@
These metrics are complementary, not redundant. Plasticity Score asks “did the agent remain adaptable on the data it collected?”, while $\Delta(\mathrm{GAP}_{\text{e}})$ asks “did that adaptation carry to perturbed tests?”. Our activation designs primarily target sustained train-side plasticity (strict floor, moderate leak, $C^{1}$ transition)—a prerequisite for learning under shift. We report both metrics to surface, not mask, this open question.
Activation HalfCheetah-v5 Humanoid-v5 Ant-v5 Hopper-v5 $\Delta_{IQM}$
Swish (SiLU) 533.55  $\pm$ 1314.41 0.00  $\pm$ 0.00 780.79  $\pm$ 730.11 13.39  $\pm$ 37.27 273.47
PReLU 839.60  $\pm$ 596.03 -316.28  $\pm$ 2211.72 94.17  $\pm$ 660.28 -22.35  $\pm$ 74.09 35.91
Sigmoid -288.53  $\pm$ 2576.85 18.92  $\pm$ 111.38 276.48  $\pm$ 1018.48 152.14  $\pm$ 129.02 85.53
Smooth-Leaky 847.49  $\pm$ 1611.57 0.00  $\pm$ 0.00 44.09  $\pm$ 1089.34 27.01  $\pm$ 72.58 35.55
Rand. Smooth-Leaky -49.38  $\pm$ 863.55 0.00  $\pm$ 0.00 -336.13  $\pm$ 971.75 -68.68  $\pm$ 96.31 59.03
Table 4: Change in Generalization Gap ($\Delta\mathrm{GAP}$). We report the difference in generalization gap between the final and first cycles ($\Delta(\mathrm{GAP}_{\text{e}})=\mathrm{GAP}_{\text{3,e}}-\mathrm{GAP}_{\text{1,e}}$). Positive values indicate a widening gap (worse transfer) over time. Values are mean $\pm$ 95% CI. The rightmost column reports the Interquartile Mean (IQM) of these deltas across environments. Note: Runs exhibiting physics simulation failures (reward <500<-500) are treated as having a gap of $0.0$ (indicating no meaningful performance difference to overfit).
@@ 8 Conclusion and Future Work
Acknowledgments
This material is based on work supported by the National Science Foundation under Grant No. 2218063. The authors acknowledge the Vermont Advanced Computing Center (VACC) at the University of Vermont for providing computational resources that have contributed to the research results reported in this paper.
@@ Appendix A Characterization of Activation Function Properties
ReLU Nair and Hinton (2010)
@@
GeLU Hendrycks and Gimpel (2016)
@@
*Gradients are small but non‑zero except at extreme inputs (risk of effective inactivity via saturation). ELU is $C^{1}$ only for $\alpha=1$. SELU has a small derivative jump at $x=0$ because $\alpha\neq 1$. SELU’s self-normalizing behavior holds under the prescribed $(\lambda,\alpha)$. Unconstrained activations are smooth, non-monotonic, and non-saturating; outputs can grow large under training. Rational do not have “slopes” like Leaky-ReLU/PReLU; they learn polynomial coefficients. Each branch retains the ReLU dead zone, but concatenation ensures that for any input at least one branch is active, avoiding global inactivity. SwiGLU is a gating mechanism (acting on a vector split) rather than a scalar activation $\phi(z)$. While it lacks a static hard dead zone based on input sign, the multiplicative gating creates a dynamic dead zone: if the gate branch saturates to zero, gradients for the value branch vanish. Proposed in this work. Potentially non-monotonic depending on the choice of slope parameter (and on the randomization bounds for R-Smooth-Leaky).
@@ Appendix B Experimental and Architectural Design
B.1 Limitations and implications of choosing an optimizer.
A valid question is how our results might interact with different optimizers (e.g., SGD, RMSProp) beyond Adam. Our study’s primary goal was to isolate the effect of activation functions on plasticity, which required fixing other key variables to create a controlled experiment. We intentionally selected the Adam optimizer for two main reasons:
  • Consistency with Prior Work: Our continual supervised learning benchmarks are adapted from recent literatureKumar et al. (2023), which used Adam for their experiments. Anchoring our design to these established baselines allows for more direct comparisons.
  • A "Stress Test" for Plasticity: More importantly, adaptive optimizers like Adam have been specifically implicated in the literature as a contributing factor to plasticity loss. Prior work has suggested that Adam can struggle to update large-magnitude weights Dohare et al. (2021) and has been shown to lose its ability to learn on non-stationary tasks Lyle et al. (2023).
We reasoned that if we wanted to study how activation functions mitigate plasticity loss, we should run our experiments in a setting known to induce it. By demonstrating that our activation-design principles can sustain plasticity even when paired with an optimizer known to exacerbate the problem, we provide a more robust finding. We do not claim this is the optimal optimizer-activation pair for SOTA results, but rather that activation design is a fundamental mechanism for preserving plasticity, even under these challenging and fixed conditions.
B.2 Limitations and implications of multi-parameter activation design.
Smooth-Leaky and Randomized Smooth-Leaky expose several shape-controlling hyperparameters. This enlarges the tuning space—and thus the cost—but also grants useful control, as reflected in our benchmarks. A common response is to replace sweeps with adaptive or learnable parameters. However, as shown in Section 3, adaptive, granular learnable slopes (e.g., PReLU) underperformed in all our settings, while bounded, stochastic constraints (e.g., RReLU) worked better. In continual learning, where short-term objectives dominate, unconstrained learned parameters can drift to task-local optima that hurt performance across tasks. Attempts to automate shape learning (Bingham and Miikkulainen, 2023; Bingham et al., 2020; Manessi and Rozza, 2018) with highly expressive forms (e.g., Rational activations Delfosse et al. (2020)) also lagged behind our first-principles designs here. This suggests that we have not yet reconciled expressivity with robust automation. A promising direction is to rethink where and on what timescale hyperparameters are adapted—potentially decoupling their updates from the main training loop to better handle non-stationarity.
B.3 CReLU integration.
@@
B.4 Hyper-parameter Sweeps for Activation Functions
@@
Leaky-ReLU | Swish / GeLU $\alpha$ | $\beta$ 0.01, 0.05, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0, 1.05, 1.1, 1.2, 1.3, 1.4, 1.5, 1.6, 1.7, 1.8, 1.9, 2.0, 2.1, 2.2, 2.3, 2.4, 2.5, 2.6, 2.7, 2.8, 2.9, 3.0
ELU / CELU $\alpha$ 0.1, 0.5, 1.0, 1.5, 2.0, 2.2, 2.4, 2.6, 2.8, 3.0, 3.3, 3.5, 3.6, 3.7, 3.8, 3.9, 4.0, 4.1, 4.2, 4.3, 4.4, 4.5, 4.6, 4.7, 4.8, 4.9, 5.0 5.1, 5.2, 5.3, 5.4, 5.5
SELU $\alpha$ 1.0, 1.3, 1.673, 2.0, 2.3, 2.4, 2.6, 2.8, 3.0, 3.1, 3.3, 3.5, 3.7, 3.8, 3.9, 4.0, 4.1, 4.2, 4.3, 4.4, 4.5, 4.6, 4.7, 4.8, 4.9, 5.0 5.1, 5.2, 5.3, 5.4, 5.5
@@
B.5 i.i.d. vs Class-Incremental Continual Learning comparison Hyperparameters
@@
B.6 Understanding Class-Incremental Continual Learning Metrics
To properly understand the choice of metrics throughout this paper, we first need to define and explain some of the classical metrics used in continual learning—Final Average Accuracy (ACC), Backward Transfer (BWT), and Forward Transfer (FWT)—and relate them to loss of plasticity. Consider a learning process unfolding over N+1N{+}1 sequential phases/tasks $t\!=\!0,\dots,N$. Between consecutive phases t1t{-}1 and $t$ ($t\!\geq\!1$), the data distribution, environment dynamics, reward signal, or task objective may change, inducing non-stationarity; we denote the distribution at phase $t$ by $E_{t}$. The model enters phase $t$ with parameters $\theta_{t-1}$ and is updated for $T_{\text{phase}}$ adaptation steps444“Steps” may be epochs, mini-batches, or timesteps, depending on the domain., producing intermediate iterates $\theta_{t,k}$ ($k\!=\!1,\dots,T_{\text{phase}}$) and terminating at $\theta_{t}\!=\!\theta_{t,T_{\text{phase}}}$.
@@ B.6 Understanding Class-Incremental Continual Learning Metrics
B.6.1 Different Accuracy metrics across Continual Learning
@@ Appendix C Expanding on Case Study 1: Negative‑Slope ‘Goldilocks Zone’
C.1 Failure Modes for Slope Magnitude.
We continue investigating the primary drivers of functional plasticity loss. As provided in the main section results, the correlation between Final Average Accuracy (ACC) and dead unit fraction remains robust and significant (Pearson r=0.51r=-0.51, $p=8.2{\times}10^{-28}$), as seen in Fig. C1, right. In contrast, the correlation between ACC and final scaled gradient norms is negligible (Pearson r=0.11r=-0.11, $p=0.029$) (Fig. C1, left). Settings that maintain dead unit fraction below approximately 8% consistently achieve high ACC (27–33%), whereas dead unit rates above 25% lead to ACC collapsing to 20% or lower. This indicates that while gradient magnitude is a factor, the rise in dead units (gradient starvation) is a more direct and reliable indicator of plasticity loss than the raw magnitude of gradients, which can be small in well-adapted, flat minima or large in failing networks with high curvature. The fact that dead units do not account for all performance variance aligns with our finding in Section 3 that landscape curvature drives failure at high slopes, distinct from unit death. The results highlight that plasticity is not about maximizing any single statistic but about achieving a delicate balance, primarily tuned by the negative slope characteristics, to ensure unit survival while maintaining a benign loss landscape.
@@
C.2 Adaptive, granular slopes are useful—but need guidance to stay in-band